OpenStudy (anonymous):

Calculate the simple interest earned when P=$15,000, r=6.8%, and t=80 days using the ordinary method. Round to the nearest cent.

OpenStudy (ranga):

Interest = P * r * t / 100 Where P is Principal = 15000 r = rate of interest = 6.8 t = time in years = 80/365 Interest = 15000 * 6.8 * 80/365 / 100 = ?

OpenStudy (anonymous):

Yes I just figured this one out!!! I got 81,600. :)

OpenStudy (anonymous):

ON my own I was so excited!

OpenStudy (anonymous):

I do have more Im stumped on can you help me Im really needing the help!

OpenStudy (anonymous):

This is one : Andre De Quadros bought a $4000, 9-month certificate of deposit (CD) that would earn 9% annual simple interest. What is the value of the CD when it matures?

OpenStudy (ranga):

How did you get 81,600?

OpenStudy (anonymous):

I took 15,000(6.8)(80) rounded to the nearest cent and got 81,600. and its one of the answers too

OpenStudy (anonymous):

what no thats not what I did but I just got the answer I have been doing this for hours now sorry I have scratch paper all over the place Im trying to find it

OpenStudy (ranga):

But that is not the right answer. You can't earn $81,600 in 86 days from a $15K investment!!!

OpenStudy (anonymous):

OH crap....man I was so proud of myself for a second ok guess start me over sorry

OpenStudy (anonymous):

its in 80 days right?

OpenStudy (ranga):

Okay 80 days. You need to convert 6.8% interest rate into decimal and so you have to divide by 100. Second, in the formula the t should be in years. Here it is in days. So divide it by 365.

OpenStudy (anonymous):

When I did what you said up top I got 22.35 and none of those is my answers I am given..... ok so this is what I have done.... 1500 * .068 * 80/365 and I got 22.35...

OpenStudy (ranga):

Is $223.56 one of the answers?

OpenStudy (anonymous):

OH YES it is duh!!!! right in front of me wow would of bit me

OpenStudy (anonymous):

sorry thanks about that

OpenStudy (ranga):

you had 1500 when it should be 15000.

OpenStudy (anonymous):

Is there anyone you can help me with the one I posted above...... Yeah I saw that now :) but I did understand thankfully LOL

OpenStudy (anonymous):

Andre De Quadros bought a $4000, 9-month certificate of deposit (CD) that would earn 9% annual simple interest. What is the value of the CD when it m

OpenStudy (anonymous):

matures sorry didnt finish the word for some reason

OpenStudy (ranga):

Similar calculation like the first problem except here they are not asking for interest but the amount at maturity. That means Principal + interest. So calculate the interest just like before. P = 40000 r = 9/100 = .09 t = 9 months = 9/12 years = 3/4 years Interest = P*r*t = ? Then add the interest to the principal P to get the maturity amount.

OpenStudy (ranga):

P = 4000 above not 40000

OpenStudy (anonymous):

no t is??? im lost on that so I times it by 9 months then divide by 12 then again by 3 and ???? see lost there

OpenStudy (ranga):

No. T should be in years. They give it in months and so we divide by 12 to convert it to years. t = 9/12 years (I just simplified it to 3/4 years. they are all the same)

OpenStudy (anonymous):

OH ok so I am dividing with a fraction then

OpenStudy (ranga):

Interest = 4000 * 0.09 * 9 / 12 = ?

OpenStudy (anonymous):

ok give me just a sec k

OpenStudy (anonymous):

ok so the slash it is a fraction or is that a divide sign cause Im still not getting the answer sorry im so frustrated

OpenStudy (ranga):

Fraction, divide sign they are all the same. Fraction is numerator divided by the denominator. Just multiply using your calculator 4000 x 0.09 x 9 first and then divide by 12.

OpenStudy (anonymous):

i see how you got 3/4 though so its a fraction right....ok sorry give me a sec here

OpenStudy (anonymous):

ok I got 270??? I did exactly what you said?

OpenStudy (ranga):

This will be the interest. You need to add the Principal $4000 to the interest to get the total amount.

OpenStudy (anonymous):

OH ok 4,270

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

OH gosh your a lifesavor!!! can you help me anymore cause i have a few

OpenStudy (ranga):

I will do one or two more.

OpenStudy (anonymous):

Susan bought a 6-month $1100 certificate of deposit. At the end of 6 months, she recieved $99 simple interest. find the annual rate of simple interest paid

OpenStudy (anonymous):

Thanks for your help means alot...not taking any math classes online anymore so confusing

OpenStudy (ranga):

It is the same formula but they ask for a different thing. Interest = P*r*t Here, P = 1100 ; t = 6 months = 0.5 year ; Interest = $99. Find r. Just plug in the numbers into the formula: 99 = 1100 * r * 0.5 Solve for r

OpenStudy (anonymous):

Ok so then I tak 1100 * 0.5 get 550 then do I divide by 99??

OpenStudy (anonymous):

to find r??

OpenStudy (ranga):

yes 99/550 = r. But that will be in decimal and usually they want r in percentage. So multiply by 100.

OpenStudy (anonymous):

5.5 *100 = 555 then ???

OpenStudy (ranga):

What is 99/550?

OpenStudy (ranga):

r is not 550/99. It is 99/550

OpenStudy (ranga):

We had 99 = 1100 * r * 0.5 99 = 550 * r divide both sides by 550 99/550 = r

OpenStudy (ranga):

Hello?

OpenStudy (anonymous):

sorry computer froze

OpenStudy (anonymous):

sorry Im having to catch up

OpenStudy (anonymous):

is 0.18

OpenStudy (anonymous):

so its 18 percent??

OpenStudy (anonymous):

are you there or did it freeze on your end Im so sorry the computers are acting up

OpenStudy (anonymous):

or internet

OpenStudy (ranga):

Yes, 18%

OpenStudy (anonymous):

Awesome :) your such a big help!!!! I dont want to bother you with more but I have some more will you be on tomorrow??? or later on ??? you always help me out

OpenStudy (ranga):

Just when I have some free time I stop by. No fixed schedule. I can just do one more here. Most of these are the same formula where you plug in the given numbers and solve for the unknown variable.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The price of an item in 2004 was $1.20. Use an inflation rate of 6% to calculate its price in 2030. Round to the nearest cent.

OpenStudy (ranga):

Have you been given a compound interest formula to use? I want to use the same symbols taught to you so I don't confuse you. Write down the formula from your book / notes / online material.

OpenStudy (anonymous):

ok sounds good Im writing everything down that you are doing thats why I take a bit longer too...

OpenStudy (anonymous):

so the same formula p x r x t

OpenStudy (ranga):

No. That is a SIMPLE interest formula. Here we need to use the COMPOUND interest formula. If you look through your notes and tell me the formula they used I will use the same. So I will wait for you to type in the formula.

OpenStudy (anonymous):

and no never been given compount intrest formula...... ok let me look

OpenStudy (anonymous):

ok so I se I=Prt A=P + I ????

OpenStudy (anonymous):

still reading

OpenStudy (ranga):

That is the simple interest formula that we have been using all along. Now it is compound interest formula.

OpenStudy (anonymous):

Yeah how they explain it has me all confused which as you can tell doesnt take much I dont see and exact formula alot of examples though not helping

OpenStudy (ranga):

It will look like this: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

Ok I see ....I didnt see it on the book.... I cant stand this book its online makes no since

OpenStudy (ranga):

That is the formula. A is the maturity amount P is the Principal r is the compounding interest in decimals n is the compounding period t is the number of years Here P = 1.2 r = 6% = 0.06 n = 1 because inflation rate is compounded once a year. t = (2030 - 2004) = 26 years Plug it into the formula and calculate A. Here you have to use exponents on your calculator.

OpenStudy (ranga):

A = 1.2(1 + 0.06)^26 (The ^ symbol means exponent) A = 1.2(1.06)^26 A = ?

OpenStudy (anonymous):

ok sorry writing it all down give me a sec :)

OpenStudy (anonymous):

33.072 is what I got

OpenStudy (ranga):

I am getting $5.46. Is that one of the answers?

OpenStudy (anonymous):

Yes but how?

OpenStudy (anonymous):

what did I do wrong?

OpenStudy (ranga):

1.06 has to be raised to 26 first using your calculator. Then multiply the result by 1.2

OpenStudy (anonymous):

OH I didnt do that!!!!!!! 1.06 what yes I did I did 1.06 * 26 then I got 27.56 then i * it by 1.2 was that wrong

OpenStudy (ranga):

26 is an exponent. So 1.06 raised to 26

OpenStudy (anonymous):

so you dont times it? sorry I know prob dumb question but need to ask

OpenStudy (ranga):

No they are completely different. 10^4 means 10 * 10 * 10 * 10 = 10000 (10)(4) means 10 * 4 = 40 completely different results.

OpenStudy (anonymous):

OH OK makes since now......

OpenStudy (anonymous):

Wow your so good at math wish I could have your knowledge!!!! I appreciate your time as always I have like 5 more Im going to try to do on my own wish me luck LOL

OpenStudy (anonymous):

And have another paper which makes no since to no one around here...I asked you the other day and no one was able to help on that one I think Im stuck on that lol

OpenStudy (ranga):

That was probably some simple concept but the terminology and the notation that they used was not something I have come across before. But if I had the class notes or text or online material I could quickly go over it and I should be able to answer those questions.

OpenStudy (anonymous):

Well thank you so very much and when your on here I hate to bug you but you explain it where I understand so Im sorry If I bug you ..... I hope your on here tomorrow for that other paper is so so very hard! I think I can do this with what you have showed me tonight with this one though. Thank you so very much!!!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Night :)